What is the energy of a photon with a frequency of 7.00 x 104 s 1? (h 6.626x 10 34J.s) J
Answer #1

We have the formula for energy of photon

E = hc / Invalid Equation

Where

h = Planks constant = 6.626 \times 10-34 J.s

c = speed of light = 3.0 \times 108 m/s

Invalid Equation = wavelength of radiation

Lets calculate Invalid Equation with given frequency

we know that

wavelength Invalid Equation = speed of light (c) / Frequency (v)

Frequency = 7.0 \times 1014 s-1

then

wavelength Invalid Equation =   3.0 \times 108 m/s / 7.0 \times 1014 s-1 = 4.29 \times 10-7 m

Now

E = hc / Invalid Equation

= 6.626 \times 10-34 J.s  \times 3.0 \times 108 m/s / 4.29 \times 10-7 m

= 4.63 \times 10-19 J

So our answer is 4.63 \times 10-19 J